# Scalable Training of Trustworthy and Energy-Efficient Predictive Graph Foundation Models for Atomistic Materials Modeling: A Case Study with HydraGNN

### v3

@@ IV-D Scalable HPO
The hyperparameter tuning has spanned important architectural hyperparameters described in Table IV. The range of architectural hyperparameters covers regions of the hyperparameter space that allow to construct HydraGNN models of extremely diverse size, which include the SMALL model and the LARGE models described in Table III as extremes.
@@
TABLE IV: Set of architectural HydraGNN hyperparameters tuned by scalable HPO.
@@
We ran four consecutive HPO runs of progressively increasing scale, each restarting from the output of the previous one. During each one of the four HPO runs, each HPO trial is associated with an independent ‘srun’ execution of the SLURM scheduler and occupies 128 Frontier nodes (i.e., 1,024 AMD Instinct MI250x GCDs) for distributed training using DDP. Concurrent HPO trials are executed asynchronously, and the termination of an HPO trial is immediately followed by the start of a new one on the same set of compute nodes. Throughout the entire execution of each HPO run, we used the open-source Omnistat data-collection infrastructure highlighted in Section IV-A to simultaneously collect GPU telemetry with measurements saved on a local Lustre PFS. Since each HPO trial is submitted as a separate job step within the global job submission, Omnistat post-processing allows us to isolate power and energy measurements at the job-step level thus providing an aggregate GPU energy consumed estimate for each trial independently. In order to ensure that the HPO process is performed in an energy-efficient way on OLCF-Frontier, we early stop the training of HydraGNN models for each HPO trial after 10 epochs. This number of epochs allows to early stop the HPO trials that are clearly underperforming in a timely manner, without wasteful energy consumption caused by further training epochs that would not likely improve their accuracy, while still ensuring that promising HPO trials are distinguishable and selected for the next computational tasks. This approach results into impactful energy savings. Our use of DeepHyper for asynchronous Bayesian optimization, combined with in-band telemetry monitoring and a strategic deployment of early termination strategies, showcases a significant advancement in the field, optimizing GNN training in ways that have not been documented prior to this work.
@@
The first HPO run used 2,048 Frontier nodes in parallel, thereby allowing 16 distinct HydraGNN architectures to be concurrently trained. This first HPO run allowed us to perform a first uninformed exploration of the hyperparameter space, and construct some guidance for the consecutive HPO runs. The results of the first HPO run have been used as inputs to inform the second HPO run, which used 3,072 Frontier nodes in parallel, thereby allowing 28 distinct HydraGNN architectures to be concurrently trained. The results of the second HPO run have been used as input to inform the third HPO run, which use 4,096 Frontier nodes in parallel, thereby allowing 32 distinct HydraGNN architectures to be concurrently trained. Similarly, the output of the third HPO run has been used to successfully guide the fourth and last HPO run, which used 8,560 Frontier nodes in parallel, thereby allowing 67 distinct HydraGNN architectures to be concurrently trained. This extensive scale of the last HPO run not only tests the limits of scalability and efficiency in computational resources, but also addresses the challenges associated with the high dimensionality of the hyperparameter space that needs to be explored, ensuring that a proper balance between exploitation and exploration is maintained.
@@
Figures 15, 16, 17, and 18 show the validation mean absolute error (MAE) as a function of wall-clock time for each one of the four HPO runs. For each one of the four HPO runs, the scattered distribution of blue dots (corresponding to values of the validation MAE for different HPO trials) shows that the HPO maintains a good degree of exploration throughout the entire execution. The red solid line indicates the minimum validation MAE obtained at a given time during the HPO run. The fact that the red line progressively lowers as time progresses confirms that HPO progressively identifies GFM architectures with better accuracy. Moreover, we notice that the HPO trials of the last run are highly concentrated around the minimum (red line) across consecutive HPO runs, thereby confirming that previous HPO runs effectively provided meaningful information to hone in narrow regions of the hyperparameter space where accurate HydraGNN models can be found.
@@
IV-E Energy Profiling
To quantify the energy usage as a function of different model sizes, three training epochs of the SMALL, MEDIUM, and LARGE model configurations listed in Table III were completed with the Omnistat telemetry tool sampling at one second intervals. Measurements consider the entire run duration including I/O for the initial data loading process. Each model executed using 1,024 GPUs on 128 compute nodes which is the minimum node count needed to accommodate memory requirements for the LARGE model configuration. The resulting GPU energy measurements as a function of model size are summarized in Table V. While total energy scales with the execution time, note that GPU utilization (occupancy) also influences the energy consumed. Table V includes mean utilization observed across all 1,024 GPUs. The LARGE case showed the highest GPU utilization–around 89%. The underling power histories used to compute total energy consumed for each model configuration are shown in Fig. 20. From these plots, we see evidence of the underlying training process with three epoch cycles visible in the power response. Furthermore, the increased GPU utilization for the larger models leads to increased GPU power demand with the LARGE model encountering peak power measurements in excess of 520 watts (W) (the peak TDP power for the AMD MI250 socket is 560W).
@@ IV-E Energy Profiling
TABLE V: Energy usage during training on OLCF-Frontier.
@@
We used this analysis to obtain energy-efficiency in the development and pretraining of GFMs as described in the following Section IV-F.
@@
IV-F Energy-efficient full training of best performing HydraGNN models identified by scalable HPO
@@
To ensure trustworthiness of our GFM, we refined the pretraining of an ensemble of twenty HPO trials selected from the four consecutive HPO runs and use them for the ensemble UQ. To this effect, we sorted the HPO trials for increasing values of the validation MAE. The results showed that nine HPO trials were outperforming all the other in terms of accuracy, with a validation MAE below 0.11. The HydraGNN architectures associated with these HPO trials have been identified as worth being further trained and thus included in the ensemble for UQ as first tier. For the selection of the remaining eleven HPO trials to be included in the ensemble, we noticed that thirty-two HPO trials were clustered within a second tier with a narrow range of the validation MAE between 0.12 and 0.125. Although these models were very close to each other in terms of accuracy, they were quite different in terms of energy consumed during the HPO run. Therefore, to ensure energy efficiency, we decided to apply a second screening of the HPO trials based on energy consumption. To this end, we performed local sorting restricted only among these thirty-two HPO trials based on increasing amount of energy consumed. The eleven models with the lowest energy consumption have been selected and included in the ensemble for UQ as second tier. In Figure 22 we show the scatterplot of the energy consumption against the validation MAE for each HPO trial executed during the four consecutive HPO runs. The nine HPO trials belonging to the first tier (in red) sit in the left most part of the plot with lowest validation MAE, whereas the eleven HPO trials of the second tier (in pink) sit along the Pareto front of energy consumption vs. validation MAE. The plot also clearly shows that all the HPO trials selected (first and second tier) sit on the side of the Pareto front that partially favors preserving optimal accuracy over energy efficiency.
@@
All twenty selected HPO trials selected use the EGNN as MPNN layer. Among the different types of MPNN layers tested, the EGNN is an equivariant model. Since equivariant models are supposed to be more data-efficient than non-equivariant models by taking full advantage of symmetries to achieve the maximal accuracy with the minimal requirement on the data, the fact that the automated HPO favors EGNN seems reasonable.
@@
The twenty selected HydraGNN models have been fully trained for a maximum of 40 epochs to reach convergence of the training, with the option of early stopping if validation MAE does not decrease across ten consecutive epochs. We report the trend of the training loss for all twenty models in Fig. 23. The training loss flattens at the end of the training history, indicating that the models have reached their maximum predictive capacity. We then used the ensemble of the fully pre-trained twenty GFMs to provide ensemble predictions with epistemic uncertainty measurement, as described in the following Section IV-G.
@@
Figure 23: Full training of twenty selected models from HPO, displaying training and test losses. Nine models are shown for the first tier and eleven models are shown for the second tier.
IV-G Epistemic UQ on ensemble of pre-trained GFMs
Figure 24: Parity plot of ensemble predictions (using twenty pretrained models from Section IV-F) of energy and atomic forces for 20,480 atomistic structures from the test split in pretraining. The red diagonal line represents the perfect model prediction.
Figure 25: Histogram of epistemic uncertainty in energy and force predictions for each dataset used in pretraining.
UQ is essential for the trustworthiness of GFMs. We calculate the uncertainties of the ensemble predictions in pretraining using the twenty models from Section IV-F. Fig. 24 presents the parity plot of ensemble averaged predictions for 20,480 atomistic structures from the test split, showing that our ensemble models have learned the generic information from the diverse pretraining data. While the accuracy is not as high as that reported in task-specific AI work, our pre-trained models are expected to reduce the fine-tuning efforts across a broad set of domain-specific tasks. Moreover, the ensemble averaging further stabilized the predictions of the GFMs.
Figure 25 shows the histogram of epistemic uncertainty of the corresponding energy and force predictions in pretraining, calculated on 4,096 atomistic structure from the test split of each dataset (in total, 20,480 structures). As expected, datasets qm7x and ANI1x, consisting of 9 million organic compounds, exhibit higher uncertainties than the other datasets, which include more than 145 million inorganic compounds. Some of the atomistic structures from the datasets qm7x and ANI1x are associated with significantly inaccurate predictions of atomic forces, which correspond to the data samples sitting on the horizontal point cloud in the parity plot for the prediction of atomic forces. Since the ensemble of GFMs correctly associate high uncertainty to atomistic structures whose predictions of atomic forces are notably inaccurate, this corroborates the fact that our GFMs are self-aware of their own regime of trustworthiness, which is an important indicator of robustness.
@@ IV-G Epistemic UQ on ensemble of pre-trained GFMs
V Conclusions and Future Work
@@
We used distributed data management capabilities to partition large volumes of data across distributed computing resources and efficiently exchange data samples across devices using low-latency communication methods. This helped preserve global data shuffling, which is crucial for maintaining good convergence of the GFM training. By scaling HPO on over 91% of the exascale OLCF-Frontier supercomputer, we have assessed the importance of thoroughly exploring a large set of hyperparameter configurations to identify HydraGNN architectures with high predictive accuracy. The use of Omnistat tools at extreme scale allowed us to conduct a thorough analysis to select HydraGNN architectures with high predictive accuracy and low energy cost. Moreover, access to exceptionally performing large scale computing facilities allowed us to develop and test ensemble UQ capabilities to measure the degree of confidence associated with the HydraGNN predictions, which contributes to achieving trustworthiness. Performing HPO and ensemble UQ at unprecedented scale on supercomputing facilities confirms our computational readiness in using HydraGNN to develop trustworthy (i.e, accurate and reliable) and energy-efficient GFMs to support the US-DOE materials science needs by providing robust and transferable computational capabilities for AI-accelerated materials discovery and design.
Future work will be devoted to improving the scaling performance of the GFM pretraining on thousands of GPUs by exploring and developing more sophisticated strategies to mitigate load imbalance between GPUs, while maintaining high quality in the attainable trustworthiness of the pre-trained GFMs. We also plan to apply techniques to read the metadata in parallel to improve the overall bandwidth of reading large-volume of datasets on supercomputers. Once the GFMs are pre-trained, we will deploy them to downstream tasks for fine-tuning, where we will illustrate the efficacy of our GFMs in reducing the amount of training data and computational resources needed to develop robust and transferable DL models for domain-specific applications.

### v4

@@ 4.4 Scalable HPO
The hyperparameter tuning has spanned important architectural hyperparameters described in Table 4. The range of architectural hyperparameters covers regions of the hyperparameter space that allow to construct HydraGNN models of extremely diverse size, which include the SMALL model and the LARGE models described in Table 3 as extremes.
@@
Table 4: Set of architectural HydraGNN hyperparameters tuned by scalable HPO.
@@
We ran four consecutive HPO runs of progressively increasing scale, each restarting from the output of the previous one. During each one of the four HPO runs, each HPO trial is associated with an independent ‘srun’ execution of the SLURM scheduler and occupies 128 Frontier nodes (i.e., 1,024 AMD Instinct MI250x GCDs) for distributed training using DDP. Concurrent HPO trials are executed asynchronously, and the termination of an HPO trial is immediately followed by the start of a new one on the same set of compute nodes. Throughout the entire execution of each HPO run, we used the open-source Omnistat data-collection infrastructure highlighted in Section 4.1 to simultaneously collect GPU telemetry with measurements saved on a local Lustre PFS. Since each HPO trial is submitted as a separate job step within the global job submission, Omnistat post-processing allows us to isolate power and energy measurements at the job-step level thus providing an aggregate GPU energy consumed estimate for each trial independently. In order to ensure that the HPO process is performed in an energy-efficient way on OLCF-Frontier, we early stop the training of HydraGNN models for each HPO trial after 10 epochs. This number of epochs allows to early stop the HPO trials that are clearly underperforming in a timely manner, without wasteful energy consumption caused by further training epochs that would not likely improve their accuracy, while still ensuring that promising HPO trials are distinguishable and selected for the next computational tasks. This approach results in impactful energy savings. Our use of DeepHyper for asynchronous Bayesian optimization, combined with in-band telemetry monitoring and a strategic deployment of early termination strategies, showcases a significant advancement in the field, optimizing GNN training in ways that have not been documented prior to this work.
@@
The first HPO run used 2,048 Frontier nodes in parallel, thereby allowing 16 distinct HydraGNN architectures to be concurrently trained. This first HPO run allowed us to perform a first uninformed exploration of the hyperparameter space, and construct some guidance for the consecutive HPO runs. The results of the first HPO run have been used as inputs to inform the second HPO run, which used 3,072 Frontier nodes in parallel, thereby allowing 28 distinct HydraGNN architectures to be concurrently trained. The results of the second HPO run have been used as input to inform the third HPO run, which use 4,096 Frontier nodes in parallel, thereby allowing 32 distinct HydraGNN architectures to be concurrently trained. Similarly, the output of the third HPO run has been used to successfully guide the fourth and last HPO run, which used 8,560 Frontier nodes in parallel, thereby allowing 67 distinct HydraGNN architectures to be concurrently trained. This extensive scale of the last HPO run not only tests the limits of scalability and efficiency in computational resources, but also addresses the challenges associated with the high dimensionality of the hyperparameter space that needs to be explored, ensuring that a judicious balance between exploitation and exploration is maintained.
@@
Figures 15, 16, 17, and 18 show the validation mean absolute error (MAE) as a function of wall-clock time for each one of the four HPO runs. For each one of the four HPO runs, the scattered distribution of blue dots (corresponding to values of the validation MAE for different HPO trials) shows that the HPO maintains a good degree of exploration throughout the entire execution. The red solid line indicates the minimum validation MAE obtained at a given time during the HPO run. The fact that the red line progressively descends as time progresses confirms that HPO progressively identifies GFM architectures with better accuracy. Moreover, we notice that the HPO trials of the last run are highly concentrated around the minimum (red line) across consecutive HPO runs, thereby confirming that previous HPO runs effectively provided meaningful information to hone in narrow regions of the hyperparameter space where accurate HydraGNN models can be found.
@@
4.5 Energy profiling
To quantify the energy usage as a function of different model sizes, three training epochs of the SMALL, MEDIUM, and LARGE model configurations listed in Table 3 were completed with the Omnistat telemetry tool sampling at one second intervals. Measurements consider the entire run duration including I/O for the initial data loading process. Each model executed using 1,024 GPUs on 128 compute nodes which is the minimum node count needed to accommodate memory requirements for the LARGE model configuration. The resulting GPU energy measurements as a function of model size are summarized in Table 5. While total energy scales with the execution time, note that GPU utilization (occupancy) also influences the energy consumed. Table 5 includes mean utilization observed across all 1,024 GPUs. The LARGE case showed the highest GPU utilization–around 89%. The underling power histories used to compute total energy consumed for each model configuration are shown in Fig. 20. From these plots, we see evidence of the underlying training process with three epoch cycles visible in the power response. Furthermore, the increased GPU utilization for the larger models leads to increased GPU power demand with the LARGE model encountering peak power measurements in excess of 520 watts (W) (the peak TDP power for the AMD MI250 socket is 560W).
@@ 4.5 Energy profiling
Table 5: Energy usage during training on OLCF-Frontier.
@@
We used this analysis to obtain energy-efficiency in the development and pre-training of GFMs as described in the following Section 4.6.
@@
4.6 Energy-efficient full training of best performing HydraGNN models identified by scalable HPO
@@
To ensure trustworthiness of our GFM, we refined the pretraining of an ensemble of fifteen HPO trials selected from the four consecutive HPO runs and use them for the ensemble UQ. To this end, we sorted the HPO trials for increasing values of the validation MAE. The results showed that four HPO trials were outperforming all the other in terms of accuracy, with a validation MAE below 0.10. The HydraGNN architectures associated with these HPO trials have been identified as worth being further trained and thus included in the ensemble for UQ as first tier. For the selection of the remaining eleven HPO trials to be included in the ensemble, we noticed that thirty-two HPO trials were clustered within a second tier with a narrow range of the validation MAE between 0.10 and 0.125. Although these models were very close to each other in terms of accuracy, they were quite different in terms of energy consumed during the HPO run. Therefore, to ensure energy efficiency, we applied a second screening of the HPO trials based on energy consumption. To this end, we performed local sorting restricted only among these thirty-two HPO trials based on increasing amount of energy consumed. The eleven models with the lowest energy consumption have been selected and included in the ensemble for UQ as second tier. In Figure 22 we show the scatterplot of the energy consumption against the validation MAE for each HPO trial executed during the four consecutive HPO runs. The four HPO trials belonging to the first tier (in red) sit in the left most part of the plot with lowest validation MAE, whereas the eleven HPO trials of the second tier (in pink) sit along the Pareto front of energy consumption vs. validation MAE. The plot also clearly shows that all the HPO trials selected (first and second tier) sit on the side of the Pareto front that partially favors preserving optimal accuracy over energy efficiency.
@@
All fifteen selected HPO trials selected use the EGNN as MPNN layer. Among the different types of MPNN layers tested, the EGNN is an equivariant model. Since equivariant models are supposed to be more data-efficient than non-equivariant models by taking full advantage of symmetries to achieve the maximal accuracy with the minimal requirement on the data, the fact that the automated HPO favors EGNN seems reasonable.
@@
The fifteen selected HydraGNN models have been fully trained for a maximum of 30 epochs to reach convergence of the training, with the option of early stopping if validation MAE does not decrease across ten consecutive epochs and with the option of saving the last performed epochs if the wall-clock allocation time is about to expire. We report the trend of the training loss for all fifteen models in Fig. 23. The training loss flattens at the end of the training history, indicating that the models have reached their maximum predictive capacity. Moreover, the models seem to reach similar final accuracy, thereby confirming that the HPO has indeed thoroughly spanned the hyperparameter space and identified HydraGNN models with similar predictive performance. We then used the ensemble of the fully pre-trained fifteen GFMs to provide ensemble predictions with epistemic uncertainty measurement, as described in the following Section 4.7.
@@
Figure 23: Full training of fifteen selected models from HPO, displaying training and test losses. Four models are shown for the first tier and eleven models are shown for the second tier.
4.7 Ensemble predictions and epistemic UQ on ensemble of pre-trained GFMs
UQ is essential for the trustworthiness of GFMs. We calculate the uncertainties of the ensemble predictions in pretraining using the fifteen models from Section 4.6. Table 6 provides the MAE and the root mean-squared error (RMSE) of the ensemble predictions of energy and forces on the training-validation-testing splits of each dataset. Given the limited number of training epochs used for the pre-training of the GFM ensemble, the accuracy achieved is very promising and indicative that the GFMs are indeed learning the underlying chemical principles that describe interactions between atoms of different constituents. Fig. 24 presents the parity plot of ensemble averaged predictions for 20,480 atomistic structures from the test split, showing that our ensemble models have learned the generic information from the diverse pretraining data. While the accuracy is not as high as that reported in task-specific AI work, our pre-trained GFMs are expected to stabilize zero-shot predictions via ensemble averaging and reduce the fine-tuning efforts across a broad set of domain-specific tasks.
Dataset Split Energy (MAE) eV/atom Energy (RMSE) eV/atom Forces (MAE) eV/angstrom Forces (RMSE) eV/angstrom
ANI1x Train 0.001742 0.003885 0.012941 0.034096
Validation 0.001640 0.002754 0.012350 0.027084
Test 0.001660 0.003009 0.012782 0.031976
MPTrj Train 0.260078 0.407861 0.172559 0.848933
Validation 0.255583 0.389681 0.149686 0.831007
Test 0.248065 0.376361 0.146531 0.638788
OC2020 Train 0.017441 0.025182 0.095165 0.183093
Validation 0.017758 0.026273 0.094872 0.185488
Test 0.021786 0.032004 0.106405 0.219948
OC2022 Train 0.045176 0.378741 0.082496 0.334222
Validation 0.049585 0.079223 0.079433 0.312630
Test 0.065060 0.144687 0.082543 0.314317
qm7x Train 0.013291 0.020044 0.141055 0.226754
Validation 0.013056 0.020049 0.140915 0.233489
Test 0.013865 0.021830 0.143753 0.235468
Table 6: MAE and RMSE values for energy and forces across datasets and splits.
Figure 25 shows the histogram of epistemic uncertainty (left column) of the corresponding energy and force predictions in pretraining, calculated on 4,096 atomistic structure from the test split of each dataset (in total, 20,480 structures). The uncertainty $\sigma_{\tilde{y}}$ is defined as the STD of ensemble predictions in Equation (6), which is commonly used but biased toward large true values and may not accurately reflect actual prediction quality. We also plot the relative uncertainty to data STD (right column) in Figure 25. Three datasets are observed to higher relative uncertainties—ANI1x, qm7x, and MPTrj—due to two factors. First, datasets ANI1x and qm7x each contain approximately 4-5 million organic compounds, in contrast to the other datasets, which include more than 145 million inorganic compounds. Second, dataset MPTrj includes higher fidelity solutions but is much smaller in size compared to the others. Some of the atomistic structures from the datasets qm7x and ANI1x are associated with significantly inaccurate predictions of atomic forces, which correspond to the data samples sitting on the horizontal point cloud in the parity plot for the prediction of atomic forces. Since the ensemble of GFMs correctly associate high uncertainty to atomistic structures whose predictions of atomic forces are notably inaccurate, this corroborates the fact that our GFMs are self-aware of their own regime of trustworthiness, which is an important indicator of robustness.
@@ 4.7 Ensemble predictions and epistemic UQ on ensemble of pre-trained GFMs
Figure 24: Parity plot of ensemble predictions (using fifteen pretrained models from Section 4.6) of energy and atomic forces for 20,480 atomistic structures from the test split in pretraining. The red diagonal line represents the perfect model prediction. We present the parity plots for each dataset separately in the supplementary material.
Figure 25: Histogram of epistemic uncertainty $\sigma_{\tilde{y}}$ (left column) and relative uncertainty $\sigma_{\tilde{y}}/\sigma_{D}$ (right column) (with $\sigma_{D}$ being the data STD for each dataset and being summarized in the supplementary material) in energy and force predictions for each dataset used in pretraining.
5 Conclusions and Future Work
@@
We used distributed data management capabilities to partition large volumes of data across distributed computing resources and efficiently exchange data samples across devices using low-latency communication methods. This helped preserve global data shuffling, which is crucial for maintaining good convergence of the GFM training. By scaling HPO on over 91% of the exascale OLCF-Frontier supercomputer, we have assessed the importance of thoroughly exploring a large set of hyperparameter configurations to identify HydraGNN architectures with high predictive accuracy. The use of Omnistat tools at extreme scale allowed us to conduct a thorough analysis to select HydraGNN architectures with high predictive accuracy and low energy cost. Moreover, access to exceptionally performing large scale computing facilities allowed us to develop and test ensemble UQ capabilities to measure the degree of confidence associated with the HydraGNN predictions, which contributes to achieving trustworthiness. Performing HPO and ensemble UQ at unprecedented scale on supercomputing facilities confirms our computational readiness in using HydraGNN to develop trustworthy and energy-efficient GFMs to support the US-DOE materials science needs by providing robust and transferable computational capabilities for AI-accelerated materials discovery and design.
Future work will be devoted to improving the scaling performance of the GFM pre-training on thousands of GPUs by exploring and developing more sophisticated strategies to mitigate load imbalance between GPUs, while maintaining high quality in the attainable trustworthiness of the pre-trained GFMs. We also plan to apply techniques to read the metadata in parallel to improve the overall bandwidth of reading large-volume of datasets on supercomputers. Once the GFMs are pre-trained, we will deploy them to downstream tasks for fine-tuning, where we will illustrate the efficacy of our GFMs in reducing the amount of training data and computational resources needed to develop robust and transferable DL models for domain-specific applications.
Code availability
The code used to develop, pre-train, and use the pre-trained GFMs is available at the following branch https://github.com/ORNL/HydraGNN/tree/Predictive_GFM_2024 of the ORNL GitHub repository for HydraGNN. More specifically, the scripts used to pre-process the data, generate the ADIOS files, run HPO, and continue the pre-training are available in the directory called examples/multidataset_hpo. The scripts that use the pre-trained ensemble of GFMs for ensemble averaging and epistemic UQ are available in the directory called examples/ensemble_learning.
Data availability
The set of ADIOS files that contain the pre-processed data and the ensemble of pre-trained GFMs that can be used for zero-shot inference and fine-tuning on downstream tasks are available open-source on the OLCF Data Constellation Facility. Release currently under review at OLCF with DOI 10.13139/OLCF/2474799.